# DINER: Debiasing Aspect-based Sentiment Analysis with Multi-variable Causal Inference

### v1

@@ 2.2 Causal Inference-based Debiasing
Causal inference Pearl (1995, 2009) has been widely employed for debiasing in various fields, including computer vision, recommendation, and natural language processing Niu et al. (2021); Zhang et al. (2021b); Tian et al. (2022). The main methods employed consist of counterfactual reasoning and causal intervention. Niu et al. (2021) proposed to remove the language bias in vision question answering by subtracting the results of a counterfactual language-only model from the results of a vanilla language-vision model. Following this work, counterfactual reasoning is widely applied to debiasing the spurious correlation between input and label in tasks including natural language understanding Tian et al. (2022), machine reading comprehension Guo et al. (2023); Zhu et al. (2023) and fact verification Xu et al. (2023). Liu et al. (2022b) proposed to de-confound the object from the context in object detection with backdoor adjustment, where an inverse probability weight approximation is made to estimate the do operator. Another way to estimate the do operator is known as normalized weighted geometrical mean (NWGM), which is firstly adopted in image caption by Liu et al. (2022a). Following this line of work, backdoor adjustment-based debiasing has widely been explored in tasks including named entity recognition Zhang et al. (2021a) and multi-modal fake news detection Chen et al. (2023). Some methods also employ other causal inference techniques, including instrument variable Wang et al. (2022) and colliding effects Zheng et al. (2022). However, most of the present debiasing methods focus on debiasing a single input variable, while we are the first to debias two input variables in ABSA simultaneously.
@@ 3.1 Structural Causal Model of ABSA
where $T\!I\!E_{a,r}$ denotes the Total Indirect Effect ($T\!I\!E$) from $A$ and $R$ on $L$, $T\!E_{a,r}$ denotes the Total Effect ($T\!I\!E$), $N\!D\!E$ denotes the Natural Direct Effect ($N\!D\!E$), and $I\!E_{a,r}$ denotes the Interaction Effect ($I\!E$) between $A$ and $R$. The total effect $T\!E$ contains all causal effects from $A$ and $R$ on $L$, inducing the biases, while the natural direct effect ($N\!D\!E$) only measures the direct causal effect between two variables, which can be regarded as the bias-only effect. Therefore, subtracting $N\!D\!E_{a}$ and $N\!D\!E_{r}$ from $T\!E_{a,r}$ will results in the unbiased causal effect from $A$ and $R$ on $L$, which is the total indirect effect $T\!I\!E_{a,r}$. It is worth noticing that since there is no causality between $A$ and $R$, the value of the interaction effect $I\!E_{a,r}$ can be set to 0.
@@
$T\!E_{a,r}$ $=L_{\boldsymbol{a},\boldsymbol{r},\boldsymbol{k}}-L_{\boldsymbol{a}^{*},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}$ (2)
$N\!D\!E_{a}$ $=L_{\boldsymbol{a},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}-L_{\boldsymbol{a}^{*},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}$ (3)
$N\!D\!E_{r}$ $=L_{\boldsymbol{a},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}-L_{\boldsymbol{a}^{*},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}$ (4)
where $L$ denotes the prediction and $\boldsymbol{x^{*}}$ denotes variable $\boldsymbol{x}$ is set to be void, we can have:
$\begin{split}T\!I\!E_{a,r}&=L_{\boldsymbol{a},\boldsymbol{r},\boldsymbol{k}}-L_{\boldsymbol{a}^{*},\boldsymbol{r},\boldsymbol{k}^{*}}-L_{\boldsymbol{a},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}+L_{\boldsymbol{a}^{*},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}\\ &=T\!I\!E_{a,r^{{}^{\prime}}}-N\!D\!E_{a}\end{split}$ (5)
@@ 3.2 Deconfounding the Review Branch with Backdoor Adjustment
Based on Eq. (5), we can estimate each component and obtain an unbiased prediction. However, as $R$ and $L$ are indirectly confounded with context $C$, it is not easy to calculate $L_{\boldsymbol{a},\boldsymbol{r},\boldsymbol{k}}$ and $L_{\boldsymbol{a}^{*},\boldsymbol{r},\boldsymbol{k}^{*}}$. Therefore, we debias the review $R$ first.
$L_{\boldsymbol{a},\boldsymbol{r},\boldsymbol{k}}=\Psi(\zeta_{\boldsymbol{a}},\zeta_{\boldsymbol{r}^{{}^{\prime}}},\zeta_{\boldsymbol{k}})$ (6)
where $\zeta_{\boldsymbol{k}}$ denotes the logit of the softmax layer, $\Psi(\cdot)$ denotes the fusion function, specially $\zeta_{\boldsymbol{r}^{{}^{\prime}}}$ denotes the debiased output based on $R$.
Consider the SCM only contains $R$, $C$, and $L$, $C$ satisfies the backdoor criterion, and we can have:
@@
where the $do(R)$ operator denotes a causal intervention that severs the direct effect of $R$ on $L$. A common workaround is the application of the Normalized Weighted Geometric Mean (NWGM) Xu et al. (2015) to approximate the effects of the $do$-operator Our approach diverges from previous methodologies by adopting an Inverse Probability Weighting (IPW) Pearl (2009) perspective, which provides a novel lens through which to approximate the infinite sampling of $(i,\boldsymbol{r})|\boldsymbol{c}$ as shown in Eq. (7).
In a finite dataset, the observable instances of $(i,\boldsymbol{r})$ for each unique $\boldsymbol{c}$ are limited. Consequently, the number of $\boldsymbol{c}$ values considered in our equation equates to the available data samples rather than the theoretically infinite possibilities of $\boldsymbol{c}$. Backdoor adjustment bridges the gap between the confounded and de-confounded models, allowing us to treat samples from the confounded model as if they were drawn from the de-confounded scenario. This leads to an approximation:
$P(L|do(R=\boldsymbol{r}))\approx\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}(L,R=\boldsymbol{r}^{k}|C),$ (8)
where $\widetilde{P}$ denotes the inverse weighted probability. For simplicity, subsequent discussions will omit $C=\boldsymbol{c}$, though it is understood that $\boldsymbol{r}$ remains dependent on $\boldsymbol{c}$. We employ a multi-head strategy, inspired by Vaswani et al. (2017) to refine the granularity of our sampling by partitioning the weight and feature dimensions into $K$ groups.
@@
$\begin{split}\widetilde{P}(L=i,R=\boldsymbol{r}^{k})\propto E(i,\boldsymbol{r}^{k};\boldsymbol{w}_{i}^{k})\\ =\tau\frac{f(i,\boldsymbol{r}^{k};\boldsymbol{w}_{i}^{k})}{g(i,\boldsymbol{r}^{k};\boldsymbol{w}_{i}^{k})},\end{split}$ (9)
@@
with $\tau$ serving as a scaling factor analogous to the inverse temperature in Gibbs distributions Geman and Geman (1984). The function $f(i,\boldsymbol{r}^{k};\boldsymbol{w}_{i}^{k})$ represents the unnormalized effect, calculated as logits, while $g(i,\boldsymbol{r}^{k};\boldsymbol{w}_{i}^{k})$ serves as a normalization term(or propensity score Austin (2011)), ensuring balance in variable magnitudes.
The computation of logits for $P(L|do(R=\boldsymbol{r}))$ is thus expressed as:
$P(L|do(R))$ $=\frac{\tau}{K}\sum_{k=1}^{K}\frac{(\boldsymbol{w}_{i}^{k})^{\top}(\ddot{\boldsymbol{r}}^{k}+\boldsymbol{d}^{k})}{(\lVert\boldsymbol{w}_{i}^{k}\rVert+\epsilon)\lVert\boldsymbol{r}^{k}\rVert}$
$=\frac{\tau}{K}\sum_{k=1}^{K}\frac{(\boldsymbol{w}_{i}^{k})^{\top}\boldsymbol{r}^{k}}{(\lVert\boldsymbol{w}_{i}^{k}\rVert+\epsilon)\lVert\boldsymbol{r}^{k}\rVert}.$ (10)
@@
Now we need to obtain context features given the current samples to do a hybrid intervention to force the model to concentrate on the debiased review. We assume $U$ as a confounder set $\{u_{i}\}^{K}_{i=1}$, where $K$ is the number of aspects in dataset and $u_{i}$ is the prototype for the context of class $i$ in feature space. Review features can be linearly or non-linearly represented by the manifolds Turk and Pentland (1991); Candès et al. (2011), and so are the context features. Therefore, we model the review-specific context features $C$ of current samples as follows:
$C=f(\boldsymbol{r},U)=\sum_{k=1}^{K}P(u_{k}|\boldsymbol{r})u_{k}$ (11)
where $P(u_{i}|\boldsymbol{r})$ is the classification probability of the feature $r$ belonging to the context of class $i$.
@@
The last remaining difficulty is implementing the contextual confounder set $U$. To obtain more useful contextual information, we employ the lower $\mathcal{K}$ layers of the model on $R\rightarrow L$ branch, which is in early training to model $U$. It is motivated by three primary considerations: First, the acknowledgment of the intrinsic wealth of contextual semantic information harbored within pre-trained language models  Liu et al. (2019b); Devlin et al. (2019) due to their extensive pre-training. Second, our requirement is not for highly advanced semantics but rather for contextual information Zeiler and Fergus (2014); Liu et al. (2022b); previous empirical studies Jawahar et al. (2019); Liu et al. (2019a); Geva et al. (2021) have shown that encoder-only models exhibit superior performance in capturing contextual information at lower layers. Third, in the initial stages of training, the model’s classification capabilities predominantly rely on context. To be specific, we encode each $\boldsymbol{r}$ using the aforementioned method, and if $\boldsymbol{r}$ contains a specific aspect, it is then represented as the representation of the corresponding $u_{k}$, and we apply the mean feature as the final $u_{k}$ representation.
@@
Given the modeling of $U$ and $C$, we are ready for the representations of context bias. We model them as $\boldsymbol{r}_{c}=\mathcal{F}(\boldsymbol{r},C)$. Following Liu et al. (2022b), we choose $W\cdot concat(r,M)$ to map since adding more networks to learn how much we need from the context is better.
@@
Now we can do Counterfactual Reasoning to debias the impact of $C$ on $R$  ($C\rightarrow R$). The final definition of debiased $\boldsymbol{r}^{{}^{\prime}}$ is as follows:
@@
$\zeta_{\boldsymbol{r}^{{}^{\prime}}}=\frac{\tau}{K}\sum_{k=1}^{K}\frac{(\boldsymbol{w}_{i}^{k})^{\top}}{(\lVert\boldsymbol{w}_{i}^{k}\rVert+\gamma)}\left(\frac{\boldsymbol{r}^{k}}{\lVert\boldsymbol{r}^{k}\rVert}-\frac{\boldsymbol{r}^{k}_{c}}{\lVert\boldsymbol{r}^{k}_{c}\rVert}\right)$ (12)
@@ 3.3 Decorrelating the Aspect Branch with Counterfactual Reasoning
While we have successfully mitigated contextual bias in the $R\rightarrow L$ pathway, the ABSA model, as delineated in Figure 2, remains susceptible to aspect-only bias. This bias persists because the prediction, denoted as $L_{\boldsymbol{a},\boldsymbol{r^{{}^{\prime}}},\boldsymbol{k}}$, is directly influenced by the aspect variable $A$. To address this, we introduce a counterfactual reasoning approach that estimates the direct causal effect of $A$ on $L$, effectively isolating the influence of $R$ and $K$. Figure 3 shows the causal graph of the counterfactual world for ABSA which describes the scenario when $A$ is set to different values $\boldsymbol{a}$ and $\boldsymbol{a}^{*}$. We also set $R$ to its reference value $\boldsymbol{r}^{*}$, therefore $K$ would attain the value $\boldsymbol{k}^{*}$ when $R=\boldsymbol{r}^{*}$ and $A=\boldsymbol{a}^{*}$. In this way,the inputs of $R$ and $K$ are blocked, and the model can only rely on the given aspect $\boldsymbol{a}$ for detection. The natural direct effect ($N\!D\!E$) of $A$ on $L$, which represents the aspect-only bias, is calculated as follows:
$N\!D\!E_{a}=L_{\boldsymbol{a},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}-L_{\boldsymbol{a}^{*},\boldsymbol{r}^{*},\boldsymbol{k}^{*}}$ (13)
@@
$L_{a,r,k}$ $=L(A=\boldsymbol{a},R=\boldsymbol{r^{{}^{\prime}}},K=\boldsymbol{k})$
$=\Psi(\zeta_{\boldsymbol{a}},\zeta_{\boldsymbol{r^{{}^{\prime}}}},\zeta_{\boldsymbol{k}})$
$=\zeta_{\boldsymbol{k}}+\tanh(\zeta_{\boldsymbol{a}})+\tanh(\zeta_{\boldsymbol{r^{{}^{\prime}}}})$ (14)
where $\zeta_{\boldsymbol{r^{{}^{\prime}}}}$ is the output of the review-only branch (i.e., $R\rightarrow L$ ), $\zeta_{\boldsymbol{a}}$ is the output of the aspect-only branch (i.e., $A\rightarrow L$ ), and $\zeta_{\boldsymbol{k}}$ is the output of fused features branch (i.e., $K\rightarrow L$ ) as shown in Figure 3. $T\!I\!E$ is the debiased result we used for inference.

### v2

@@ 2.2 Causal Inference-based Debiasing
Causal inference Pearl (1995, 2009) has been widely employed for debiasing in various fields, including computer vision, recommendation, and natural language processing Niu et al. (2021); Zhang et al. (2021b); Tian et al. (2022). The main methods employed consist of counterfactual reasoning and causal intervention. Niu et al. (2021) proposed to remove the language bias in vision question answering by subtracting the results of a counterfactual language-only model from the results of a vanilla language-vision model. Following this work, counterfactual reasoning is widely applied to debiasing the spurious correlation between input and label in tasks including natural language understanding Tian et al. (2022), machine reading comprehension Guo et al. (2023); Zhu et al. (2023) and fact verification Xu et al. (2023). Liu et al. (2022b) proposed to de-confound the object from the context in object detection with backdoor adjustment, where an inverse probability weight approximation is made to estimate the do-operator. Another way to estimate the do-operator is known as normalized weighted geometrical mean (NWGM), which is firstly adopted in image caption by Liu et al. (2022a). Following this line of work, backdoor adjustment-based debiasing has widely been explored in tasks including named entity recognition Zhang et al. (2021a) and multi-modal fake news detection Chen et al. (2023). Some methods also employ other causal inference techniques, including instrument variable Wang et al. (2022) and colliding effects Zheng et al. (2022). However, most of the present debiasing methods focus on debiasing a single input variable, while we are the first to debias two input variables in ABSA simultaneously.
@@ 3.1 Structural Causal Model of ABSA
where $T\!I\!E_{a,r}$ denotes the Total Indirect Effect ($T\!I\!E$) from $A$ and $R$ on $L$, $T\!E_{a,r}$ denotes the Total Effect ($T\!E$), $N\!D\!E$ denotes the Natural Direct Effect ($N\!D\!E$), and $I\!E_{a,r}$ denotes the Interaction Effect ($I\!E$) between $A$ and $R$. The total effect $T\!E$ contains all causal effects from $A$ and $R$ on $L$, inducing the biases, while the natural direct effect ($N\!D\!E$) only measures the direct causal effect between two variables, which can be regarded as the bias-only effect. Therefore, subtracting $N\!D\!E_{a}$ and $N\!D\!E_{r}$ from $T\!E_{a,r}$ will results in the unbiased causal effect from $A$ and $R$ on $L$, which is the total indirect effect $T\!I\!E_{a,r}$. It is worth noticing that since there is no causality between $A$ and $R$, the value of the interaction effect $I\!E_{a,r}$ can be set to 0.
@@
$T\!E_{a,r}$ $=L_{a,r,k}-L_{a^{*},r^{*},k^{*}}$ (2)
$N\!D\!E_{a}$ $=L_{a,r^{*},k^{*}}-L_{a^{*},r^{*},k^{*}}$ (3)
$N\!D\!E_{r}$ $=L_{a^{*},r,k^{*}}-L_{a^{*},r^{*},k^{*}}$ (4)
where $L$ denotes the prediction and $x^{*}$ denotes variable $x$ is set to be void, we can have:
$\begin{split}T\!I\!E_{a,r}&=L_{a,r,k}-L_{a^{*},r,k^{*}}-L_{a,r^{*},k^{*}}+L_{a^{*},r^{*},k^{*}}\\ &=T\!I\!E_{a,r^{{}^{\prime}}}-N\!D\!E_{a}\end{split}$ (5)
where $r^{{}^{\prime}}$ denotes the debiased review, obtained after the process of deconfounding.
@@ 3.2 Deconfounding the Review Branch with Backdoor Adjustment
Based on Eq. (5), we can estimate each component and obtain an unbiased prediction. However, as $R$ and $L$ are indirectly confounded with context $C$, it is not easy to calculate $L_{a,r,k}$ and $L_{a^{*},r,k^{*}}$. Therefore, we debias the review $R$ first.
$L_{a,r,k}=\Psi(\zeta_{a},\zeta_{r^{{}^{\prime}}},\zeta_{k})$ (6)
where $\zeta_{k}$ denotes the logit of the softmax layer, $\Psi(\cdot)$ denotes the fusion function, specially $\zeta_{r^{{}^{\prime}}}$ denotes the debiased output based on $R$.
There are mainly three types of causal intervention methods based on causal inference, the backdoor adjustment, the front-door adjustment, and the instrument variable adjustment. However, the front-door adjustment requires a mediator variable between input and output, which is not applicable in our SCM Zhang et al. (2024a, b). The instrument variable adjustment involves building up an extra instrument variable in SCM, which makes the already complex SCM even more complex Wang et al. (2022). So we choose the backdoor adjustment for debiasing the review branch. Consider the SCM only contains $R$, $C$, and $L$, $C$ satisfies the backdoor criterion, and we can have:
@@
where the $do(R)$ operator denotes a causal intervention that severs the direct effect of $R$ on $L$.
A common workaround is the application of the Normalized Weighted Geometric Mean (NWGM) Xu et al. (2015) to approximate the effects of the $do$-operator. Our approach adopts an Inverse Probability Weighting (IPW) Pearl (2009) perspective, which provides a novel lens through which to approximate the infinite sampling of $(l,r)|c$ as shown in Eq. (7).
In a finite dataset, the observable instances of $(l,r)$ for each unique $c$ are limited. Consequently, the number of $c$ values considered in our equation equates to the available data samples rather than the theoretically infinite possibilities of $c$. Backdoor adjustment bridges the gap between the confounded and de-confounded models, allowing us to treat samples from the confounded model as if they were drawn from the de-confounded scenario. This leads to an approximation:
$\begin{split}P(L|do(R=r))&\approx\widetilde{P}(L,R|C=c)\\ &\approx\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}(L,R=r^{k}|C=c)\end{split}$ (8)
where $\widetilde{P}$ denotes the inverse weighted probability. We employ a multi-head strategy, inspired by Vaswani et al. (2017) to refine the granularity of our sampling by partitioning the weight and feature dimensions into $K$ groups, $r_{k}$ denotes the review information in the group $K=k$. For simplicity, subsequent discussions will omit $C=c$, though it is understood that $r$ remains dependent on $c$. We will employ $T\!D\!E$ to debias this effect following Liu et al. (2022a) and Tang et al. (2020).
@@
$\begin{split}\widetilde{P}(L=l,R=r^{k})\propto E(l,r^{k};w^{k})\\ =\tau\frac{f(l,r^{k};w^{k})}{g(l,r^{k};w^{k})}\end{split}$ (9)
@@
with $\tau$ serving as a scaling factor analogous to the inverse temperature in Gibbs distributions Geman and Geman (1984), $w^{k}$ denotes the weight parameter in the group $K=k$. The numerator $f(l,r^{k};w^{k})$ represents the unnormalized effect, calculated as logits $(w^{k})^{\top}r^{k}$, while the denominator $g(l,r^{k};w^{k})$ serves as a normalization term(or propensity score Austin (2011)), which ensures balanced magnitudes of the variables. The denominator, i.e., inverse probability weight, becomes the propensity score under the energy-based model, where the effect is divided into the controlled group $\lVert w^{k}\rVert\cdot\lVert r^{k}\rVert$ and the uncontrolled group $\epsilon\cdot\lVert r^{k}\rVert$.
The computation of logits for $P(L|do(R=r))$ is thus expressed as:
$P(L|do(R))$ $=\frac{\tau}{K}\sum_{k=1}^{K}\frac{(w^{k})^{\top}r^{k}}{(\lVert w^{k}\rVert+\epsilon)\lVert r^{k}\rVert}$ (10)
@@
Now we need to obtain context features given the current samples to force the model to concentrate on the debiased review based on $T\!D\!E$. We assume $U$ as a confounder set $\{u_{i}\}^{N}_{i=1}$, where $N$ is the number of aspects in dataset and $u_{i}$ is the prototype for the context of class $i$ in feature space. Review features can be linearly or non-linearly represented by the manifolds Turk and Pentland (1991); Candès et al. (2011), and so are the context features. Therefore, we model the review-specific context features $C$ of current samples as follows:
$C=f(r,U)=\sum_{N=1}^{N}P(u_{n}|r)u_{n}$ (11)
where $P(u_{i}|r)$ is the classification probability of the feature $r$ belonging to the context of class $i$.
@@
The last remaining difficulty is implementing the contextual confounder set $U$. To obtain more useful contextual information, we employ the lower $\mathcal{K}$ layers of the model on $R\rightarrow L$ branch, which is in early training to model $U$. It is motivated by three primary considerations: First, the acknowledgment of the intrinsic wealth of contextual semantic information harbored within pre-trained language models  Liu et al. (2019b); Devlin et al. (2019) due to their extensive pre-training. Second, our requirement is not for highly advanced semantics but rather for contextual information Zeiler and Fergus (2014); Liu et al. (2022b); previous empirical studies Jawahar et al. (2019); Liu et al. (2019a); Geva et al. (2021) have shown that encoder-only models exhibit superior performance in capturing contextual information at lower layers. Third, in the initial stages of training, the model’s classification capabilities predominantly rely on context.
To be specific, we encode each $r$ using the aforementioned method, and if $r$ contains a specific aspect, it is then represented as the representation of the corresponding $u_{n}$, and we apply the mean feature as the final $u_{n}$ representation.
@@
Given the modeling of $U$ and $C$, we are ready for the representations of context bias. We model them as $r_{c}=\mathcal{F}(r,C)$. Following Liu et al. (2022b), we choose $W\cdot concat(r,M)$ to map since adding more networks to learn how much we need from the context is better.
@@
Now we can debias the impact of $C$ on $R$  ($C\rightarrow R$) based on $T\!D\!E$. The final definition of debiased $r^{{}^{\prime}}$ is as follows:
@@
$\zeta_{r^{{}^{\prime}}}=\frac{\tau}{K}\sum_{k=1}^{K}\frac{(w^{k})^{\top}}{(\lVert w^{k}\rVert+\epsilon)}\left(\frac{r^{k}}{\lVert r^{k}\rVert}-\frac{r^{k}_{c}}{\lVert r^{k}_{c}\rVert}\right)$ (12)
@@ 3.3 Decorrelating the Aspect Branch with Counterfactual Reasoning
While we have successfully mitigated contextual bias in the $R\rightarrow L$ pathway, the ABSA model, as delineated in Figure 2, remains susceptible to aspect-only bias. This bias persists because the prediction, denoted as $L_{a,r^{{}^{\prime}},k}$, is directly influenced by the aspect variable $A$. To address this, we introduce a counterfactual reasoning approach that estimates the direct causal effect of $A$ on $L$, effectively isolating the influence of $R$ and $K$. Figure 3 shows the causal graph of the counterfactual world for ABSA which describes the scenario when $A$ is set to different values $a$ and $a^{*}$. We also set $R$ to its reference value $r^{*}$, therefore $K$ would attain the value $k^{*}$ when $R=r^{*}$ and $A=a^{*}$. In this way,the inputs of $R$ and $K$ are blocked, and the model can only rely on the given aspect $a$ for detection. The natural direct effect ($N\!D\!E$) of $A$ on $L$, which represents the aspect-only bias, is calculated as follows:
$N\!D\!E_{a}=L_{a,r^{*},k^{*}}-L_{a^{*},r^{*},k^{*}}$ (13)
@@
$L_{a,r,k}$ $=L(A=a,R=r^{{}^{\prime}},K=k)$
$=\Psi(\zeta_{a},\zeta_{r^{{}^{\prime}}},\zeta_{k})$
$=\zeta_{k}+\tanh(\zeta_{a})+\tanh(\zeta_{r^{{}^{\prime}}})$ (14)
where $\zeta_{r^{{}^{\prime}}}$ is the output of the review-only branch (i.e., $R\rightarrow L$ ), $\zeta_{a}$ is the output of the aspect-only branch (i.e., $A\rightarrow L$ ), and $\zeta_{k}$ is the output of fused features branch (i.e., $K\rightarrow L$ ) as shown in Figure 3. $T\!I\!E$ is the debiased result we used for inference.